What is the mass of 4.91 moles of gold
1
Expert's answer
2020-05-12T11:35:07-0400

If you keep in mind pure 100% gold, then

n=m(Ag)/Mr(Ag) =>

=> m(Ag)=n*Mr(Ag)=4,91*197=967,27g